OpenStudy (anonymous):

lim x approaches zero cosx^2-1/x

OpenStudy (anonymous):

which is -sin(x) / x, which is -1

OpenStudy (anonymous):

this is \[\frac{cos^2(x)-1}{x}=\frac{-\sin^2(x)}{x}\]

OpenStudy (anonymous):

since lim x-->0 of sinx / x is 1

OpenStudy (anonymous):

oh no not quite

OpenStudy (anonymous):

wow im off today my bad i was rushing!

OpenStudy (anonymous):

hopitals rule

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\cos^2(x)-1}{x}=\lim_{x \rightarrow 0}\frac{\cos(x)-1}{x} \cdot \lim_{x \rightarrow 0}(\cos(x)+1)=0 \cdot (\cos(0)+1)=0(1+1)=0(2)=0\]

OpenStudy (anonymous):

-2sin(x) / 1 so lim is zero

OpenStudy (anonymous):

same as \[-\frac{\sin(x)}{x}\times \sin(x)\] right? first limit is 1second is 0 get 0

OpenStudy (anonymous):

satellite isnt first limit -1... im pretty sure im right for that

OpenStudy (anonymous):

no need for l'hopital here. limit is straight up 0

OpenStudy (anonymous):

yeah it is -1 you are right. product is 0 in any case for sure

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well u can do either product of lims or hopital...wtvr u want

OpenStudy (anonymous):

thanks